Microeconomics multiple choice questions

If implicit costs equal accounting profit, economic profit must be

Select one:

a. negative.

b. positive.

c. zero.

d. higher in the short run than in the long run.

e. higher in the long run than in the short run.

The reason the change in total cost divided by the change in output is equal to the change in total variable cost divided by the change in output, is because

Select one:

a. total variable cost rises as output rises.

b. of the law of diminishing marginal returns.

c. total fixed cost does not change as output changes.

d. total cost does not change as output changes.

The main difference between the short run and the long run is that

Select one:

a. firms earn losses in the long run, but not in the short run.

b. the long run always refers to a time period of one year or longer.

c. in the long run, only one input can be fixed.

d. in the short run, one or more inputs are fixed.

e. none of the above
Answer #1

1 . If implicit costs equal accounting profit, economic profit must be zero.

Accounting Profits = Total Revenue- Explicit Costs ..................1

Economic Profits = Total Revenue - Explicit Costs - Implicit Costs....................2

Economic Profits = Accounting Profit - Implicit Cost (From 1)

Since accounting profits and implicit costs are the same (Given in the question)

Economic Profits = Accounting Profit - Accounting Profits

= 0

2.The reason the change in total cost divided by the change in output is equal to the change in total variable cost divided by the change in output, is because total fixed cost does not change as output changes.

Total fixed cost remains constant irrespective of output. You would have to pay the same fixed cost irrespective of whether you produce one good or 100 goods. The only change which occurs is change in Variable Cost.

We Know that Total Cost = Total Fixed Cost + Total Variable Cost

Consider TC1 = TFC + TVC1 ( Say cost of producing one units)...............1

TC2 = TFC + TVC2 (Say Cost of Producing two units)...............2

IF we subtract 1 from 2 we get

TC2 - TC1 = TFC + TVC2 -TFC - TVC1

  TC2 - TC1 = TVC2 - TVC1

3. The main difference between the short run and long run is that in the short run one or more inputs are fixed.

Reason: Long Run is a time period where all the factors of production are variable. (Thus option "c" is wrong.)

The long run is not time specific. It is industry specific. For eg: For an ice cream seller his long run would be one month or one year while for a steel industry , it's long run would be greater than five years. (Hence option "b" is wrong)

Long run curve is also called a planning curve. We as humans never plan losses . Thus it is expected that firms would earn profits in the long run (Thus optio "a" is wrong)

Thus only option "d" is closest to the answer.